OpenStudy (day_2010):

Can someone help? I thought I solved it correct and got X=-6. The problem is: solve. 0.04(4x-6)=0.16(x+7)-1.36 The answer was supposed to be "all real numbers". How?

OpenStudy (phi):

All values for x will work in your equation. Did you distribute the 0.04 to get 0.16x - 0.24 and distribute the 0.16 to get 0.16x+1.12-1.36 so your equation is 0.16x - 0.24 = 0.16x+1.12-1.36 or 0.16x - 0.24 = 0.16x-0.24 which is always true, for any value of x

OpenStudy (day_2010):

I assumed that you would change the decimals to whole numbers... does it work that way too?

OpenStudy (day_2010):

I think I was trying to solve for x some how...

OpenStudy (phi):

You could multiply both sides by 100 to get 16x-24= 16x-24 If you continue to simplify, by subtracting 16x from both sides and add 24 to both sides 0=0 when you get an equation like this, it means the original will always be true, for all values of x

OpenStudy (day_2010):

I think I was trying to solve for x and that is how I got -6.... I put all the x's to one side..

OpenStudy (day_2010):

Thanks for clarifying! :) much appreciated!

OpenStudy (phi):

I think I was trying to solve for x and that is how I got -6.... I put all the x's to one side.. If you did this you should see the x cancel. So you must have made a mistake. You should be able to get to 0.16x - 0.24 = 0.16x-0.24 now if you move the x from the right side by add -0.16x to both sides: 0.16x - 0.24 -0.16x= 0.16x-0.24 -0.16x all the x's disappear -0.24= -0.24

OpenStudy (day_2010):

I don't know who I did that, funny.... but I do understand what your saying. Thanks for the help. :)